THE COF_{2} HOT BANDS IN THE ν_{6} REGION AND THE 2ν_{6} OVERTONE^{a}

\underline{E. A. COHEN}, B. J. DROUIN, and L. R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099.

Using results from our analysis of rotational and infrared spectra of ν_{2}, ν_{3}, ν_{5}, and ν_{6}^{b} we have fit the hot bands in the ν_{6} region near 774 cm^{−1} originating in ν_{2}, ν_{3}, and ν_{5} as well as the 2ν_{6} overtone band. Although the 2ν_{6} state is perturbed by the nearby and as yet unassigned ν_{2} + ν_{3} and ν_{2} + ν_{5} states, a set of effective parameters was derived which adequately describes the 2ν_{6} spectrum as well as the 2ν_{6} − ν_{6} hot band. These results improve the description of the ν_{6} spectrum, particularly in the Q-branch region. The molecular parameters will be presented and examples of spectra will be shown.

aThis work was done at the Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration.

bE.A. Cohen, and B.J. Drouin, Abstract TE10, International Symposium on Molecular Spectroscopy, Columbus, OH, 2010.